In 1865, Jules Verne proposed sending men to the Moon by firing a space capsuile from a 220-m-long cannon with fral speed of 10.97 km/s. What would have been the stand an acceleration of 15g for a short time.) m/근 large acceieration experenced by she spece travelers during their launch? (A human an Compare your answer with the free-fall acceleration, 9.80 m/s2 Need Help? es IWatch N An object moves with constant acceleration 3.60 m/s? and over a time interval reaches a final velocity of 10.2 m/s. (a) If its original velocity is 5.10 m/s, what is its displacement during the time interval? (b) What is the distance it travels during this interval? c) If its initial velocity is +5 10 m/s, what is its displacement during this interval? (d) What is the total distance it travels during the linterval in part (c)? Need Hetp? ilithe Submd Rights Reparved Copyright o20gge Leaming, Irig Show All FYEREadingC..docx
Answer #1

1. Vf =10.97 km/s

vf=10970 m/s

Vi=0 m/s

We know,

Vf^2 - Vi^2 = 2 a S

10970^2- 0 = 2 X a x 220

a = 273502 m/s^2

Which is 273502/9.8 = 27880 g